The Force between two identical charge's 1.90 x10^-2 C is separated by 1 cm. What is the electric froce between the two of charges?

1
Expert's answer
2022-03-06T18:07:03-0500

Force

"F=\frac{kq_1q_2}{d^2}"


"F=\frac{9\times10^9\times1.90\times10^{-2}\times1.90\times10^{-2}}{0.01^2}=3.25\times10^{10}N"